OpenStudy (anonymous):

\[ \lim_{x~\rightarrow ~\infty} \frac{ \sin x }{ 2x^2 + x } \] How to find algebraically?

OpenStudy (anonymous):

\[ \lim_{x \rightarrow \infty} \frac{ 1 }{ 2x^2+ x } * \lim_ {x \rightarrow \infty} \sin x\] ?

OpenStudy (zarkon):

squeeze theorem

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

got it from there lol thanks

OpenStudy (zarkon):

good